What is x^2 - 4x +4/ x^2 -3x +2 reduced to lowest terms
a) x-2/ x-1
b) x-2/x+1
c) x +2/ x-1
d) x+2/ x+1

Answer by rothauserc(4718)

x^2 - 4x +4/ x^2 -3x +2 =
(x-2)^2 / (x-2)(x-1) =
(x-2) / (x-1)
answer is a)